OpenStudy (anonymous):

PLASE HELP WILL MEDAL AND FAN!! Solve for n: 21k – 3n + 9p > 3p + 12. A. n > 2p + 7k – 4 B. n < 2p + 7k – 4 C. n > –4p – 7k + 4 D. n < –4p – 7k + 4

pooja195 (pooja195):

Start by adding 3p to both sides 21k−3n+9p+3p>3p+12+3p What is the expression we are left with?

OpenStudy (anonymous):

idk i dont get the problem at all @pooja195

pooja195 (pooja195):

Alright Let's go through this step by step :) We want p on one side corrcect? Good so to start off we need to start by moving all the p's to one side to do this we should start by canceling what we can 21k – 3n + 9p > 3p + 12. Start by subtract p fro both sides 3-3=0 9-3=?

OpenStudy (anonymous):

6

pooja195 (pooja195):

good now we have 21k−3n+6p>12 Any ideas on what the next step might be/

pooja195 (pooja195):

*?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Ok we are isolating p so we need to subtract 21k to both sides

pooja195 (pooja195):

21k-21k−3n+6p>12-21k What are we left with?

pooja195 (pooja195):

wait we are solving for n ....

pooja195 (pooja195):

im so sorry wait im gonna rewrite the steps Solve for n: 21k – 3n + 9p > 3p + 12. Subtract 9p 21k - 3n + 9p > 3p + 12 subtract 21k -3n > -6p - 21k + 12

pooja195 (pooja195):

Any ideas on what to do from here??

OpenStudy (anonymous):

subtract 6p?

pooja195 (pooja195):

hmm nope we need to isolate the n so just divide everything by -3 -3/-3=n -6/-3=? -21/-3=? 12/-3=?

OpenStudy (anonymous):

2 7 -4

pooja195 (pooja195):

Well done ^_^ now we have n>7k+2p−4 HOWEVER when we divide by a negative we FLIP the sign so it would be n<7k+2p−4

pooja195 (pooja195):

Do you understand the concept? :)

pooja195 (pooja195):

I hope i didnt confuse you too much

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

yes thanks

pooja195 (pooja195):

=)